Empirical Formula

2)



i) A compound is 40.0 % C, 6.70% H, and 53.3% O by mass. Assume that we have a 100g sample of this compound.
The molecular formula mass of this compound is 240 amu. What are the subscripts in the actual molecular formula?
Enter the subscripts for C, H and O respectively, separated by commas (e.g., 5,6,7).


(ii) Write the empirical formula of copper chloride based on the experimental data.
Express your answer as a chemical formula.

Please show each step and
explain each step with as much detail including why you took so and so step.
Thank You
Answer #1
Trial A
Volume of Copper Chloride Solution (mL) 49.6
Mass of Filter Paper (g) 0.908
Mass of filter paper with copper (g) 1.694

Trial B
Volume of Copper Chloride Solution (mL) 48.3
Mass of Filter Paper (g) 0.922
Mass of filter paper with copper (g) 1.693

Trial C
Volume of Copper Chloride Solution (mL) 42.2
Mass of Filter Paper (g) 0.919
Mass of filter paper with copper (g) 1.588

Take averages (you coulda done this for me):
46.7mL, 0.9163333g , 1.658333333 g with copper.
Take the difference of filter paper with and without copper to get 0.742g copper.

How concentrated is the solution?
answered by: nakoa page
Answer #2

Assuming we have 100g of this compound, we now have 40g C, 6.7 g H, 53.3 g O.
Divide each value by its respective amu because we want to know how many atoms of each , so:
C: 40/12= 3.33333
H: 6.7/1= 6.7
O: 53.3/16= 3.33333

You got some nasty numbers. Divide by the smallest value (which is 3.333333) to get the lowest coefficents you want.
C: 3.33333/3.3333= 1
H: 6.7 / 3.33333= 2
O:: 53.3 / 3.3333= 1

NIce. Now you have a ratio. C:H:O = 1:2:1
Calculate the amu of the smallest compound you can get (which is CH2O) so that you could use this to find the molecular formula.

Soooooo.... 12 + 2 + 16 = 30 amu forCH2O.

Your question says that the molecular compound formula is 240 amu, so divide 240 by 30. You get 8.

Normally, you would multiply 8 to your coefficients ofCH2O to get the answer in your form. (8, 16, 8)

b)

the next part is hard to answer b/c what's your experimental data?

Take your amount of copper and divide by its amu. Take your value of Chloride and divide it by its amu.

Now divide by the small number there (like in the top I used 3.333 because that's the smallest number there)

You should get whole numbersss... I'm not sure what'll happen after that, but if you have the molecular mass of your compound, you should do what Itried to explain above..

answered by: candie